OpenStudy (anonymous):

(tanθ+sinθ)/(2tanθ)=cos^2((θ)/(2)) verify the identity (tanθ+sinθ)/(2tanθ)=cos^2((θ)/(2)) verify the identity @Mathematics

OpenStudy (anonymous):

try converting to sin and cos = sinx / cos x + sinx / 2sinx/cosx = sin x + sinxcosx cos x -------------- * ---- cos x 2sin x = 1 + cos x -------- 2 = cos^2 (x/2)

OpenStudy (hoblos):